# On a pinning model in correlated Gaussian random environments

### v1

@@ 4. The Stratonovich case
$\|\mathrm{Tr}^{k}{\phi}_{m}\|^{2}_{|\mathcal{H}|^{\otimes(m-2k)}}$ (4.5)
$=$ $\int_{[0,1]^{2m}}\prod_{i=1}^{m-2k}|t_{i}-t_{i}^{\prime}|^{2H-2}\prod_{j=1}^{k}|s_{2j}-s_{2j-1}|^{2H-2}|s_{2j}^{\prime}-s_{2j-1}^{\prime}|^{2H-2}$
×$\times\quad\phi_{m}(s_{1},\cdots,s_{2k},t_{1},\cdots,t_{m-2k})\phi_{m}(s_{1}^{\prime},\cdots,s_{2k}^{\prime},t_{1}^{\prime},\cdots,t_{m-2k}^{\prime}){\mathrm{d}}\boldsymbol{s}{\mathrm{d}}\boldsymbol{s^{\prime}}{\mathrm{d}}\boldsymbol{t}{\mathrm{d}}\boldsymbol{t^{\prime}}$
$=$ $\mathrm{Tr}^{m}{\Phi}_{m}\leq\mathrm{Tr}^{m}{\phi}_{2m},$
where
$\Phi_{m}(t_{1},t_{1}^{\prime},\cdots,t_{m-2k},t_{m-2k}^{\prime},s_{1},\cdots s_{2k},s_{1}^{\prime},\cdots,s_{2k}^{\prime})$
::=:= $\phi_{m}(s_{1},\cdots,s_{2k},t_{1},\cdots,t_{m-2k})\phi_{m}(s_{1}^{\prime},\cdots,s_{2k}^{\prime},t_{1}^{\prime},\cdots,t_{m-2k}^{\prime})$
$\leq$ ${\phi}_{2m}(t_{1},t_{1}^{\prime},\cdots,t_{m-2k},t_{m-2k}^{\prime},s_{1},\cdots s_{2k},s_{1}^{\prime},\cdots,s_{2k}^{\prime}).$
Substituting (4.5) into (4.4) yields the desired inequality.
@@
$\quad\psi_{2m,2k}(t_{1},t_{1}^{\prime},\cdots,t_{m-2k},t_{m-2k}^{\prime},s_{1},\cdots s_{2k},s_{1}^{\prime},\cdots,s_{2k}^{\prime})$
$\geq\psi_{m,k}(s_{1},\cdots,s_{2k},t_{1},\cdots,t_{m-2k})\psi_{m,k}(s_{1}^{\prime},\cdots,s_{2k}^{\prime},t_{1}^{\prime},\cdots,t_{m-2k}^{\prime}),$
@@
$Z_{N}$ $=\mathbb{E}_{\tau}\bigg{[}\Big{(}\sum_{k=0}^{\infty}\frac{1}{k!}\Big{(}\sum_{n=1}^{N}h_{N}\sigma_{n}\Big{)}^{k}\Big{)}\Big{(}\sum_{m=0}^{\infty}\frac{1}{m!}\Big{(}\sum_{n=1}^{N}\beta_{N}\omega_{n}\sigma_{n}\Big{)}^{m}\Big{)}\bigg{]}$ (4.6)
$=\sum_{m=0}^{\infty}\sum_{k=0}^{\infty}\frac{1}{k!m!}\sum_{\boldsymbol{n}\in\llbracket N\rrbracket^{m+k}}\beta_{N}^{m}h_{N}^{k}\mathbb{P}_{\tau}(n_{1},\cdots,n_{m},\cdots,n_{m+k}\in\tau)\prod_{i=1}^{m}\omega_{n_{i}},$
$=:\sum_{m=0}^{\infty}\sum_{k=0}^{\infty}\frac{1}{k!m!}\mathbb{S}_{m,k}^{(N)}.$
@@
$\mathbb{S}_{m,k}^{(N)}$ $=\sum_{\boldsymbol{n}\in\llbracket N\rrbracket^{m+k}}\beta_{N}^{m}h_{N}^{k}\mathbb{P}_{\tau}(n_{1},\cdots,n_{m+k}\in\tau)\sum_{j=0}^{\lfloor\frac{m}{2}\rfloor}\sum_{\begin{subarray}{c}B\subset\llbracket m\rrbracket\\ |B|=m-2j\end{subarray}}\omega_{\boldsymbol{n}_{B}}^{\diamond}\mathbb{E}[\omega_{\boldsymbol{n}_{\llbracket m\rrbracket\backslash B}}]$
$=\sum_{j=0}^{\lfloor\frac{m}{2}\rfloor}\sum_{\begin{subarray}{c}B\subset\llbracket m\rrbracket\\ |B|=m-2j\end{subarray}}\beta_{N}^{m}h_{N}^{k}\sum_{\boldsymbol{n}\in\llbracket N\rrbracket^{m+k}}\mathbb{P}_{\tau}(n_{1},\cdots,n_{m+k}\in\tau)\omega_{\boldsymbol{n}_{B}}^{\diamond}\mathbb{E}[\omega_{\boldsymbol{n}_{\llbracket m\rrbracket\backslash B}}].$
For each $j=0,1,\cdots,\lfloor\frac{m}{2}\rfloor$, all the terms in the summation $\sum_{\begin{subarray}{c}B\subset\llbracket m\rrbracket,|B|=m-2j\end{subarray}}$ are equal by symmetry. Fix $j\in\{0,1,...,\lfloor\frac{m}{2}\rfloor\}$ and note that there are in total $\binom{m}{m-2j}$ subsets of $\llbracket m\rrbracket$ with cardinality $m-2j$. Without loss of generality, we may assume $B=\{2j+1,...,m\}$. Then, we have
@@
equal to each other for all pair partitions $V$. Note that there are in total $(2j-1)!!$! pair partitions of $\llbracket 2j\rrbracket$, and thus
$\mathbb{S}_{m,k}^{(N)}$ $=\sum_{j=0}^{\lfloor\frac{m}{2}\rfloor}\binom{m}{m-2j}(2j-1)!!\beta_{N}^{m}h_{N}^{k}\sum_{\boldsymbol{n}\in\llbracket N\rrbracket^{m+k}}\omega_{\boldsymbol{n}_{B}}^{\diamond}\prod_{l=1}^{j}\mathbb{E}[\omega_{n_{2l-1}}\omega_{n_{2l}}]\mathbb{P}_{\tau}(n_{1},\cdots,n_{m+k}\in\tau)$ (4.9)
@@
Compare (4.9) with Hu-Meyer’s formula (2.8) and note that $\mathrm{Tr}^{j}\psi_{m,k}\in|\mathcal{H}|^{\otimes(m-2j)}$ for $k=0,1,...,\lfloor\frac{m}{2}\rfloor$ by Lemma 4.1. From now on, without loss of generality, we assume that $m$ is odd and the case of even $m$ can be treated in the same manner. By Proposition 2.6, we have
@@
$:=N^{-2jH}N^{-k}\frac{L(N)^{m+k}}{(N^{\alpha-1})^{m+k}}\sum_{\boldsymbol{u}\in\llbracket N\rrbracket^{2j}}\sum_{\boldsymbol{q}\in\llbracket N\rrbracket^{m-2j}}\sum_{\boldsymbol{s}\in\llbracket N\rrbracket^{k}}\prod_{l=1}^{j}\mathbb{E}[\omega_{n_{2l-1}}\omega_{n_{2l}}]P_{\tau}(\boldsymbol{u},\boldsymbol{q},\boldsymbol{s})\prod_{i=1}^{m-2j}\mathbf{1}_{(\frac{q_{i}-1}{N},\frac{q_{i}}{N}]}(t_{i}),$
where we recall that $\mathbf{I}_{m-2j}^{(N)}$ is defined in (2.11), $P_{\tau}(\boldsymbol{u},\boldsymbol{q},\boldsymbol{s})$ is defined in (3.5), and note that
$\frac{L(N)^{m+k}}{N^{m(\alpha-1+H)+\alpha k}}=N^{-(m-2j)H}N^{-2jH}N^{-k}\frac{L(N)^{m+k}}{(N^{\alpha-1})^{m+k}}.$
For $j=0$, we have $Y_{N}=\mathbf{I}_{m}^{(N)}(f_{m,k}^{(N)}-\psi_{m,k})$, which converges to 00 in $L^{2}$ as $N\to\infty$ as has been proved in the proof of Proposition 3.1. For $j=1,...,\frac{m-1}{2}$, by Lemma 2.4 and Lemma A.4, we get
@@
By a similar argument used in the proof of Lemma 4.1, we can show that the right-hand side of (4.16) belongs to $L^{1/H}$. This is provides the desired $L^{1/H}$ domination.
@@ 4.3. $L^{1}$-bound of the rescaled partition functions
In this subsection, we aim to prove the uniform $L^{1}$-boundedness, under Assumption (A3), for the partition function $Z_{N}$. Recall its Taylor expansion (4.6) and by Cauchy-Schwartz inequality, we have that
$\begin{split}\mathbb{E}[Z_{N}]\leq\sum\limits_{k=0}^{\infty}\sum\limits_{m=0}^{\infty}\frac{1}{k!}\frac{1}{m!}\mathbb{E}_{\tau}\Big{[}\Big{(}\sum\limits_{n=1}^{N}h_{N}\sigma_{n}\Big{)}^{2k}\Big{]}^{\frac{1}{2}}\mathbb{E}\Big{[}\Big{(}\sum\limits_{n=1}^{N}\beta_{N}\omega_{n}\sigma_{n}\Big{)}^{2m}\Big{]}^{\frac{1}{2}}.\end{split}$ (4.17)
@@
$\sum\limits_{k=0}^{\infty}\frac{1}{k!}\mathbb{E}_{\tau}\Big{[}\Big{(}\sum\limits_{k=0}^{\infty}h_{N}\sigma_{N}\Big{)}^{2k}\Big{]}^{\frac{1}{2}}\leq\sum\limits_{k=0}^{\infty}\frac{C^{k}}{k!}\sqrt{\frac{(2k)!}{\Gamma(2k\alpha^{\prime})}}\lesssim\sum\limits_{k=0}^{\infty}\frac{C^{k}}{k^{k\alpha^{\prime}}}<\infty.$ (4.18)
@@
$\begin{split}\mathbb{E}\Big{[}\Big{(}\sum\limits_{n=1}^{N}\beta_{N}\omega_{n}\sigma_{n}\Big{)}^{2m}\Big{]}^{\frac{1}{2}}=\Big{(}\sum\limits_{\boldsymbol{n}\in\llbracket N\rrbracket^{2m}}\beta_{N}^{2m}{\mathbb{P}}_{\tau}(n_{1},\cdots,n_{2m}\in\tau)\mathbb{E}_{\omega}[\omega_{\boldsymbol{n}_{\llbracket m\rrbracket}}]\Big{)}^{\frac{1}{2}}.\end{split}$ (4.19)
Similar to the proof of Proposition 4.5, $\mathbb{E}[\omega_{\boldsymbol{n}_{\llbracket m\rrbracket}}]=\sum_{V}\prod_{\{l_{1},l_{2}\}\in V}\mathbb{E}[\omega_{l_{1}}\omega_{l_{2}}]$, where $V$ is all pair partitions of $\llbracket 2m\rrbracket$. Note that $|V|=(2m-1)!!$!. By (1.4) and (3.11), a Riemann sum approximation and the kernel modification procedure in the proof of Lemma 4.1, (4.19) is bounded above by
@@
$\sum\limits_{m=0}^{\infty}\frac{1}{m!}\mathbb{E}\Big{[}\Big{(}\sum\limits_{n=1}^{N}\beta_{N}\omega_{n}\sigma_{n}\Big{)}^{2m}\Big{]}^{\frac{1}{2}}\leq\sum\limits_{m=0}^{\infty}\frac{C^{m}}{m^{m(\alpha^{\prime}+H-\frac{3}{2})}}<\infty.$ (4.21)
@@
$\lim_{M\to\infty}\sup_{N\in\mathbb{N}}\sum_{l=M}^{\infty}\sum_{m+k=l}\frac{1}{k!m!}\mathbb{E}\Big{[}\Big{|}h_{N}\sum_{n=1}^{N}\sigma_{n}\Big{|}^{k}\Big{|}\beta_{N}\sum_{n=1}^{N}\omega_{n}\sigma_{n}\Big{|}^{m}\Big{]}=0.$ (4.22)
@@ Appendix A Some miscellaneous results
$\omega_{\boldsymbol{n}_{\llbracket m\rrbracket}}=\sum_{j=0}^{[\frac{m}{2}]}\sum_{\begin{subarray}{c}B\subset\llbracket m\rrbracket\\ |B|=m-2j\end{subarray}}\omega_{\boldsymbol{n}_{B}}^{\diamond}\mathbb{E}[\omega_{\boldsymbol{n}_{\llbracket m\rrbracket\backslash B}}],$
@@
The following Hardy-Littlewood-Sobolev inequality was proved in [14].
@@
For $H\in(\frac{1}{2},1]$, the following inequality holds:
@@
where $C_{H}$ is a constant depending only on $H$.
The following result can be obtained by direct computation.
@@
It is not hard to see that the first two sums are finite. For the third term, note that $(bm)^{ck}(ck)^{bm}\leq(bm)^{bm}(ck)^{ck}$. Let $d=c-a$ and we have that
@@
$=$ $\sum_{m\geq m_{0}}\sum_{k\geq k_{0}}\frac{1}{\left(C_{A}m\right)^{dk}\left(C^{\prime}_{A}k\right)^{bm}}\leq\sum_{m\geq m_{0}}\frac{1}{\left(C^{\prime}_{A}k_{0}\right)^{bm}}\sum_{k\geq k_{0}}\frac{1}{\left(C_{A}m_{0}\right)^{dk}}<\infty,$
@@
Let $X$ be an exponentially integrable random variable on $(\Omega,\mathcal{F},{\mathbb{P}})$ and $A\in\mathcal{F}$. Then we have
$\mathbb{E}\left[e^{X}-e^{X\mathbf{1}_{A}}\right]={\mathbb{P}}(A^{c})\left(\mathbb{E}\left[e^{X}\Big{|}A^{c}\right]-1\right)=\mathbb{E}\left[e^{X}\mathbf{1}_{A^{c}}\right]-{\mathbb{P}}(A^{c}).$
@@
$={\mathbb{P}}(A)\cdot 0+{\mathbb{P}}(A^{c})\mathbb{E}[f(X)-f(0)|A^{c}]$
$=\mathbb{E}[f(X)\mathbf{1}_{A^{c}}]-f(0){\mathbb{P}}(A^{c}),$
@@
By triangular inequality and Cauchy-Schwartz inequality, we get
@@
For any fixed $\varepsilon\in(0,1)$, we may find $K=K(\varepsilon)$ such that for all $k>K$ and $n\in\mathbb{N}$,
${\mathbb{P}}(Y_{n}>k)^{\frac{1}{2}}+{\mathbb{P}}(Y_{n}^{\varepsilon}>k)^{\frac{1}{2}}<\frac{\varepsilon}{\sqrt{M_{\varepsilon}}}\quad\text{and}\quad\mathbb{E}|Y^{\varepsilon}_{n}\mathbf{1}_{[Y^{\varepsilon}_{n}>k]}|<\varepsilon$
by Markov’s inequality and Cauchy-Schwartz inequality. Therefore, for all $k>K$ and $n\in\mathbb{N}$, $\mathbb{E}|Y_{n}\mathbf{1}_{[Y_{n}>k]}|<3\varepsilon$. This proves the uniform integrability of $\{Y_{n},n\in\mathbb{N}\}$. ∎
Acknowledgements. J. Song is partially supported by NSFC Tianyuan Key Program Project (No. 12226001), National Natural Science Foundation of China (No. 12471142), and the Fundamental Research Funds for the Central Universities. R. Wei is supported by Xi’an Jiaotong-Liverpool University Research Development Fund (No. RDF-23-01-024) and NSFC (No. 12401170). The authors thank Quentin Berger, Rongfeng Sun and Jinjiong Yu for reading the first version of this paper and providing invaluable suggestions.

### v2

@@ 4. The Stratonovich case
$\|\mathrm{Tr}^{k}{\phi}_{m}\|^{2}_{|\mathcal{H}|^{\otimes(m-2k)}}=$ $\int_{[0,1]^{2m}}\prod_{i=1}^{m-2k}|t_{i}-t_{i}^{\prime}|^{2H-2}\prod_{j=1}^{k}|s_{2j}-s_{2j-1}|^{2H-2}|s_{2j}^{\prime}-s_{2j-1}^{\prime}|^{2H-2}$ (4.5)
$\times\quad\phi_{m}(s_{1},\cdots,s_{2k},t_{1},\cdots,t_{m-2k})\phi_{m}(s_{1}^{\prime},\cdots,s_{2k}^{\prime},t_{1}^{\prime},\cdots,t_{m-2k}^{\prime}){\mathrm{d}}\boldsymbol{s}{\mathrm{d}}\boldsymbol{s^{\prime}}{\mathrm{d}}\boldsymbol{t}{\mathrm{d}}\boldsymbol{t^{\prime}}$
$\leq$ $\mathrm{Tr}^{m}{\phi}_{2m}.$
For the last inequality above, denoting $\boldsymbol{t}:=(t_{1},\cdots,t_{m-2k})$, $\boldsymbol{s}:=(s_{1},\cdots,s_{2k})$ and $\boldsymbol{t}^{\prime},\boldsymbol{s}^{\prime}$ correspondingly, we have that by the symmetry and the expression of $\phi$,
$\phi_{m}(\boldsymbol{s},\boldsymbol{t})\phi_{m}(\boldsymbol{s}^{\prime},\boldsymbol{t}^{\prime})\leq\phi_{2m}(\boldsymbol{s},\boldsymbol{s}^{\prime},t_{1},t^{\prime}_{1},\cdots,t_{m-2k},t^{\prime}_{m-2k}),$
since each subinterval among $(\boldsymbol{s},\boldsymbol{t})$ and $(\boldsymbol{s}^{\prime},\boldsymbol{t}^{\prime})$ is shortened by the interlacement between $(\boldsymbol{s},\boldsymbol{t})$ and $(\boldsymbol{s}^{\prime},\boldsymbol{t}^{\prime})$. Also note that the kernel $K(\cdot)$ in the trace (2.9) is not influenced by the permutation in $\phi$. Then substituting (4.5) into (4.4) yields the desired inequality.
@@
$\psi_{m,k}(\boldsymbol{s},\boldsymbol{t})\psi_{m,k}(\boldsymbol{s}^{\prime},\boldsymbol{t}^{\prime})\leq\psi_{2m,2k}(\boldsymbol{s},\boldsymbol{s}^{\prime},t_{1},t_{1}^{\prime},\cdots,t_{m-2k},t_{m-2k}^{\prime}),$
@@
$Z_{N}$ $=\mathbb{E}_{\tau}\bigg[\Big(\sum_{k=0}^{\infty}\frac{1}{k!}\Big(\sum_{n=1}^{N}h_{N}\delta_{\frac{n}{N}}\Big)^{k}\Big)\Big(\sum_{m=0}^{\infty}\frac{1}{m!}\Big(\sum_{n=1}^{N}\beta_{N}\omega_{n}\delta_{\frac{n}{N}}\Big)^{m}\Big)\bigg]$ (4.6)
$=\sum_{m=0}^{\infty}\sum_{k=0}^{\infty}\frac{1}{k!m!}\sum_{\boldsymbol{n}\in\llbracket N\rrbracket^{m+k}}\beta_{N}^{m}h_{N}^{k}\mathbb{P}_{\tau}(n_{1},\cdots,n_{m},\cdots,n_{m+k}\in\tau)\prod_{i=1}^{m}\omega_{n_{i}}=:\sum_{m=0}^{\infty}\sum_{k=0}^{\infty}\frac{1}{k!m!}\mathbb{S}_{m,k}^{(N)}.$
@@
$\mathbb{S}_{m,k}^{(N)}$ $=\sum_{\boldsymbol{n}\in\llbracket N\rrbracket^{m+k}}\beta_{N}^{m}h_{N}^{k}\mathbb{P}_{\tau}(n_{1},\cdots,n_{m+k}\in\tau)\sum_{j=0}^{\lfloor\frac{m}{2}\rfloor}\sum_{B\subset\llbracket m\rrbracket,~|B|=m-2j}\omega_{\boldsymbol{n}_{B}}^{\diamond}\mathbb{E}[\omega_{\boldsymbol{n}_{\llbracket m\rrbracket\backslash B}}]$
$=\sum_{j=0}^{\lfloor\frac{m}{2}\rfloor}\sum_{B\subset\llbracket m\rrbracket,~|B|=m-2j}\beta_{N}^{m}h_{N}^{k}\sum_{\boldsymbol{n}\in\llbracket N\rrbracket^{m+k}}\mathbb{P}_{\tau}(n_{1},\cdots,n_{m+k}\in\tau)\omega_{\boldsymbol{n}_{B}}^{\diamond}\mathbb{E}[\omega_{\boldsymbol{n}_{\llbracket m\rrbracket\backslash B}}].$
For each $j=0,1,\cdots,\lfloor\frac{m}{2}\rfloor$, all the terms in the summation $\sum_{\begin{subarray}{c}B\subset\llbracket m\rrbracket,|B|=m-2j\end{subarray}}$ are equal by symmetry. Fix $j\in\{0,1,...,\lfloor\frac{m}{2}\rfloor\}$ and note that there are in total $\binom{m}{m-2j}$ subsets of $\llbracket m\rrbracket$ with cardinality m2jm-2j. Without loss of generality, we may assume $B=\{2j+1,...,m\}$. By the Wick theorem (see, e.g., [19, Theorem 1.28]), we have
@@
equal to each other for all pair partitions $V$. Since there are in total (2j1)!!(2j-1)!! pair partitions of $\llbracket 2j\rrbracket$,
$\mathbb{S}_{m,k}^{(N)}$ $=\sum_{j=0}^{\lfloor\frac{m}{2}\rfloor}\binom{m}{m-2j}(2j-1)!\beta_{N}^{m}h_{N}^{k}\sum_{\boldsymbol{n}\in\llbracket N\rrbracket^{m+k}}\omega_{\boldsymbol{n}_{B}}^{\diamond}\prod_{l=1}^{j}\mathbb{E}[\omega_{n_{2l-1}}\omega_{n_{2l}}]\mathbb{P}_{\tau}(n_{1},\cdots,n_{m+k}\in\tau)$ (4.9)
@@
Compare (4.9) with Hu–Meyer formula (2.10) and note that $\mathrm{Tr}^{j}\psi_{m,k}\in|\mathcal{H}|^{\otimes(m-2j)}$ for $k=0,1,...,\lfloor\frac{m}{2}\rfloor$ by Lemma 4.1. From now on, without loss of generality, we assume that $m$ is odd and the case of even $m$ can be treated in the same manner. By Proposition 2.7, we have
@@
$=N^{-2jH}N^{-k}\frac{L(N)^{m+k}}{(N^{\alpha-1})^{m+k}}\!\!\sum_{\boldsymbol{u}\in\llbracket N\rrbracket^{2j}}\sum_{\boldsymbol{q}\in\llbracket N\rrbracket^{m-2j}}\sum_{\boldsymbol{s}\in\llbracket N\rrbracket^{k}}\prod_{l=1}^{j}\mathbb{E}[\omega_{n_{2l-1}}\omega_{n_{2l}}]P_{\tau}(\boldsymbol{u},\boldsymbol{q},\boldsymbol{s})\prod_{i=1}^{m-2j}\mathbf{1}_{(\frac{q_{i}-1}{N},\frac{q_{i}}{N}]}(t_{i}),$
where we recall that $\mathbf{I}_{m-2j}^{(N)}$ is defined in (2.13), $P_{\tau}(\boldsymbol{u},\boldsymbol{q},\boldsymbol{s})$ is defined in (3.5).
For $j=0$, we have $Y_{N}=\mathbf{I}_{m}^{(N)}(f_{m,k}^{(N)}-\psi_{m,k})$, which converges to 0 in $L^{2}$ as $N\to\infty$ as has been proved in the proof of Proposition 3.1. For $j=1,...,\frac{m-1}{2}$, by Lemma 2.5 and Lemma A.4, we get
@@
By a similar argument used in the proof of Lemma 4.1, we can show that the right-hand side of (4.16) belongs to $L^{1/H}$. This provides the desired $L^{1/H}$ domination.
@@ 4.3. $L^{1}$-bound of the rescaled partition functions
In this subsection, we prove the uniform $L^{1}$-boundedness, under Assumption (A3), for the partition function $Z_{N}$. Recall its Taylor expansion (4.6) and by the Cauchy-Schwarz inequality, we have that
$\begin{split}\mathbb{E}[Z_{N}]\leq\sum\limits_{k=0}^{\infty}\sum\limits_{m=0}^{\infty}\frac{1}{k!}\frac{1}{m!}\mathbb{E}_{\tau}\Big[\Big(\sum\limits_{n=1}^{N}h_{N}\delta_{\frac{n}{N}}\Big)^{2k}\Big]^{\frac{1}{2}}\mathbb{E}\Big[\Big(\sum\limits_{n=1}^{N}\beta_{N}\omega_{n}\delta_{\frac{n}{N}}\Big)^{2m}\Big]^{\frac{1}{2}}.\end{split}$ (4.17)
@@
$\sum\limits_{k=0}^{\infty}\frac{1}{k!}\mathbb{E}_{\tau}\Big[\Big(\sum\limits_{n=1}^{N}h_{N}\delta_{\frac{n}{N}}\Big)^{2k}\Big]^{\frac{1}{2}}\leq\sum\limits_{k=0}^{\infty}\frac{C^{k}}{k!}\sqrt{\frac{(2k)!}{\Gamma(2k\alpha^{\prime})}}\lesssim\sum\limits_{k=0}^{\infty}\frac{C^{k}}{k^{k\alpha^{\prime}}}<\infty.$ (4.18)
@@
$\begin{split}\mathbb{E}\Big[\Big(\sum\limits_{n=1}^{N}\beta_{N}\omega_{n}\delta_{\frac{n}{N}}\Big)^{2m}\Big]^{\frac{1}{2}}=\Big(\sum\limits_{\boldsymbol{n}\in\llbracket N\rrbracket^{2m}}\beta_{N}^{2m}{\mathbb{P}}_{\tau}(n_{1},\cdots,n_{2m}\in\tau)\mathbb{E}_{\omega}[\omega_{\boldsymbol{n}_{\llbracket m\rrbracket}}]\Big)^{\frac{1}{2}}.\end{split}$ (4.19)
Similar to the proof of Proposition 4.5, $\mathbb{E}[\omega_{\boldsymbol{n}_{\llbracket m\rrbracket}}]=\sum_{V}\prod_{\{l_{1},l_{2}\}\in V}\mathbb{E}[\omega_{l_{1}}\omega_{l_{2}}]$, where $V$ is all pair partitions of $\llbracket 2m\rrbracket$. Note that |V|=(2m1)!!|V|=(2m-1)!!. By (1.4) and (3.12), a Riemann sum approximation and the kernel modification procedure in the proof of Lemma 4.1, (4.19) is bounded above by
@@
$\sum\limits_{m=0}^{\infty}\frac{1}{m!}\mathbb{E}\Big[\Big(\sum\limits_{n=1}^{N}\beta_{N}\omega_{n}\delta_{\frac{n}{N}}\Big)^{2m}\Big]^{\frac{1}{2}}\leq\sum\limits_{m=0}^{\infty}\frac{C^{m}}{m^{m(\alpha^{\prime}+H-\frac{3}{2})}}<\infty.$ (4.21)
@@
$\lim_{M\to\infty}\sup_{N\in\mathbb{N}}\sum_{l=M}^{\infty}\sum_{m+k=l}\frac{1}{k!m!}\mathbb{E}\Big[\Big|h_{N}\sum_{n=1}^{N}\delta_{\frac{n}{N}}\Big|^{k}\Big|\beta_{N}\sum_{n=1}^{N}\omega_{n}\delta_{\frac{n}{N}}\Big|^{m}\Big]=0.$ (4.22)
@@ Appendix A Some miscellaneous results
The following two lemmas about the Wick product can be found in [19, Corollary 3.17, Theorem 3.9]. For Lemma A.1, see also [29, equation (A.5)].
@@
[…]
@@
The following lemma about the commutative diagram of weak convergence can be found in [6, Theorem 3.2, Chapter 1].
@@
The following inequality is borrowed from [23].
@@
For $H\in(\frac{1}{2},1]$, there exists a constant $C_{H}$ depending only on $H$, such that,
@@
The following result can be obtained by a direct computation.
@@
It is not hard to see that the first two sums are finite. For the third term, note that $(bm)^{ck}(ck)^{bm}\leq(bm)^{bm}(ck)^{ck}$. Let d=cad=c-a and we have that
@@
$=$ $\sum_{m\geq m_{0}}\sum_{k\geq k_{0}}\frac{1}{(C_{A}m)^{dk}(C^{\prime}_{A}k)^{bm}}\leq\sum_{m\geq m_{0}}\frac{1}{(C^{\prime}_{A}k_{0})^{bm}}\sum_{k\geq k_{0}}\frac{1}{(C_{A}m_{0})^{dk}}<\infty,$
@@
Let $X$ be an exponentially integrable random variable on $(\Omega,\mathcal{F},{\mathbb{P}})$ and $A\in\mathcal{F}$. Then
$\mathbb{E}\left[e^{X}-e^{X\mathbf{1}_{A}}\right]={\mathbb{P}}(A^{c})\Big(\mathbb{E}\Big[e^{X}\Big|A^{c}\Big]-1\Big)=\mathbb{E}\left[e^{X}\mathbf{1}_{A^{c}}\right]-{\mathbb{P}}(A^{c}).$
@@
$={\mathbb{P}}(A)\cdot 0+{\mathbb{P}}(A^{c})\mathbb{E}[f(X)-f(0)|A^{c}]=\mathbb{E}[f(X)\mathbf{1}_{A^{c}}]-f(0){\mathbb{P}}(A^{c}),$
@@
By the triangular inequality and the Cauchy-Schwarz inequality, we get
@@
For any fixed $\varepsilon\in(0,1)$, we may find $K=K(\varepsilon)$ such that for all $k>K$ and $n\in\mathbb{N}$, ${\mathbb{P}}(Y_{n}>k)^{1/2}+{\mathbb{P}}(Y_{n}^{\varepsilon}>k)^{1/2}<\varepsilon(M_{\varepsilon})^{-1/2}$ and $\mathbb{E}|Y^{\varepsilon}_{n}\mathbf{1}_{[Y^{\varepsilon}_{n}>k]}|<\varepsilon$ by Markov’s inequality and the Cauchy-Schwarz inequality. Hence, $\sup_{k\geq K,{n}\in\mathbb{N}}\mathbb{E}|Y_{n}\mathbf{1}_{[Y_{n}>k]}|<3\varepsilon$. This proves the uniform integrability of $\{Y_{n},n\in\mathbb{N}\}$. ∎
Acknowledgments
We thank Quentin Berger, Rongfeng Sun, and Jinjiong Yu for valuable suggestions on an earlier version of the paper. We also thank Julien Poisat for helpful comments and references. We are further grateful to the referees for their suggestions and comments, which have significantly improved the paper.
funding
J. Song is partially supported by NSFC (No. 12471142) and the Fundamental Research Funds for the Central Universities. R. Wei is supported by NSFC (No. 12401170) and Xi’an Jiaotong-Liverpool University Research Development Fund (No. RDF-23-01-024).